OpenStudy (anonymous):

Help! Pre-Cal What are the zeros in f(x)=x^5+x^3-6x?

OpenStudy (amistre64):

what do we get when we factor a GCF?

OpenStudy (anonymous):

Isn't the GCF 1?

OpenStudy (amistre64):

i was going for x

OpenStudy (amistre64):

f(x) = x(x^4 + x^2 -6) so we know x=0 makes this equal to 0 now to work the hidden quadratic, let a = x^2 such that a^2 + a - 6 = 0

OpenStudy (anonymous):

So x(x^4+x^2-6)?

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

Oh ok I got it now, thank you!

OpenStudy (amistre64):

youre welcome, good luck :)